You need to have $10,000 available in 25 years.If you are earning an interest rate of 5%,how much would you need to invest today to have $10,000 in 25 years, compounded annually?
Answer by cta1987(8)

A=P(1+r/n)^nt
t=25 years
annually=1 time per year
P=10000
r=0.05
A=10,000(1+0.05/1)^1*25
A=10000(1.05)^25
A=10,000(3.386354941)
A=33863.55